OpenStudy (anonymous):

The equation is x^3-12x^2+28x-9. x=2.618. I need to make this equal 0, but I can only make it equal .001. I will post my work in the comments.

OpenStudy (anonymous):

0=2.618^3-12(2.618)^2+28(2.618)-9 0=17.944-82.247+73.304-9 0=-64.303+64.304 0=.001 The answer should be .0004